### v1

# Progressive Compositionality in Text-to-Image Generative Models

Abstract
Despite the impressive text-to-image (T2I) synthesis capabilities of diffusion models, they often struggle to understand compositional relationships between objects and attributes, especially in complex settings. Existing solutions have tackled these challenges by optimizing the cross-attention mechanism or learning from the caption pairs with minimal semantic changes. However, can we generate high-quality complex contrastive images that diffusion models can directly discriminate based on visual representations? In this work, we leverage large-language models (LLMs) to compose realistic, complex scenarios and harness Visual-Question Answering (VQA) systems alongside diffusion models to automatically curate a contrastive dataset, ConPair, consisting of 15k pairs of high-quality contrastive images. These pairs feature minimal visual discrepancies and cover a wide range of attribute categories, especially complex and natural scenarios. To learn effectively from these error cases, i.e., hard negative images, we propose EvoGen, a new multi-stage curriculum for contrastive learning of diffusion models. Through extensive experiments across a wide range of compositional scenarios, we showcase the effectiveness of our proposed framework on compositional T2I benchmarks. The project page can be found at https://github.com/evansh666/EvoGen.
1 Introduction
The rapid advancement of text-to-image generative models (Saharia et al., 2022; Ramesh et al., 2022) has revolutionized the field of image synthesis, driving significant progress in various applications such as image editing (Brooks et al., 2023; Zhang et al., 2024), video generation (Brooks et al., 2024) and medical imaging (Han et al., 2024a). Despite their remarkable capabilities, state-of-the-art models such as Stable Diffusion (Rombach et al., 2022) and DALL-E 3 (Betker et al., 2023) still face challenges with composing multiple objects into a coherent scene (Huang et al., 2023; Liang et al., 2023; Majumdar et al., 2024). Common issues include incorrect attribute binding, miscounting, and flawed object relationships as shown in Figure 1. For example, when given the prompt “a red motorcycle and a yellow door”, the model might incorrectly bind the colors to the objects, resulting in a yellow motorcycle.
Recent progress focuses on optimizing the attention mechanism within diffusion models to better capture the semantic information conveyed by input text prompts (Agarwal et al., 2023; Chefer et al., 2023; Pandey et al., 2023). For example, Meral et al. (2023) proposes manipulating the attention on objects and attributes as contrastive samples during test-time to optimize model performance. While more focused, the practical application of these methods still falls short of fully addressing attribute binding and object relationships. Other works advocate to develop compositional generative models to improve compositional performance as each constituent model captures distributions of an independent domain (Du & Kaelbling, 2024). However, such approach assumes a fixed prespecified structure to compose models, limiting generalization to new distributions.
Figure 1: Limited Compositionality Understanding in Diffusion Models. Existing SOTA models such as SDXL, DALL-E 3 often fail to correctly compose objects and attributes. The bottom are images generated by our EvoGen.
In this paper, we argue that curriculum training is crucial to equip diffusion models with a fundamental understanding of compositionality. Given that existing models often struggle with even basic tasks (e.g., generating three cats when prompted with “Two cats are playing”(Wang et al., 2024), we progressively introduce more complex compositional scenarios during fine-tuning. This staged training strategy helps models build a solid foundation before tackling intricate cases before improving their performance on a wide range of compositional tasks.
Although many datasets exist for compositional generation (Wang et al., 2023; Feng et al., 2023), there remains a significant gap in datasets that offer a clear progression from simple to complex samples within natural and reasonable contexts. Moreover, creating high-quality contrastive image datasets is both costly and labor-intensive, especially given the current limitations of generative models in handling compositional tasks. To address this, we propose an automatic pipeline to generate faithful contrastive image pairs, which we find crucial for guiding models to focus on compositional discrepancies. In summary, our work can be summarized as follows:
Contrastive compositional dataset.
We introduce ConPair, a meticulously crafted compositional dataset consisting of high-quality contrastive images with minimal visual representation differences, covering a wide range of attribute categories. By leveraging LLMs, we scale up the complexity of compositional prompts while maintaining a natural context design. Our dataset features faithful images generated by diffusion models, assisted by VQA systems to ensure accurate alignment with the text prompts.
EvoGen: Curriculum contrastive learning.
We are the first work to incorporate curriculum contrastive learning into a diffusion model to improve compositional understanding. The process is broken into three streamlined sub-tasks: (1) learning single object-attribute composition, (2) mastering attribute binding between two objects, and (3) handling complex scenes with multiple objects. We conduct extensive experiments using the latest benchmarks and demonstrate that EvoGen significantly boosts the model’s compositional understanding, outperforming most baseline generative methods.
2 Preliminary Background
2.1 Diffusion models
We implement our method on top of the state-of-the-art text-to-image (T2I) model, Stable Diffusion (SD) (Rombach et al., 2022). In this framework, an encoder $\mathcal{E}$ maps a given image $x\in\mathcal{X}$ into a spatial latent code $z=\mathcal{E}(x)$, while a decoder $\mathcal{D}$ reconstructs the original image, ensuring $\mathcal{D}(\mathcal{E}(x))\approx x$.
A pre-trained denoising diffusion probabilistic model (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) for noise estimation and a pre-trained CLIP text encoder (Radford et al., 2021) to process text prompts into conditioning vectors $c(y)$. The DDPM model $\epsilon(\theta)$ is trained to minimize the difference between the added noise $\epsilon$ and the model’s estimate at each timestep $t$,
$\mathcal{L}=\mathbb{E}_{z\sim\mathcal{E}(x),y,\varepsilon\sim\mathcal{N}(0,1),t}\left[||\varepsilon-\varepsilon_{\theta}(z_{t},t,c(y))||_{2}^{2}\right].$ (1)
During inference, a latent $z_{T}$ is sampled from $\mathcal{N}(0,1)$ and is iteratively denoised to produce a latent $z_{0}$. The denoised latent $z_{0}$ is then passed to the decoder to obtain the image $x^{\prime}=\mathcal{D}(z_{0})$.
Dataset # Samples
Contra.
text
Contra.
Image
Categories Complex
DrawBench (Saharia et al., 2022) 200 3 (color, spatial, action)
CC-500 (Feng et al., 2023) 500 1 (color)
Attn-and-Exct (Chefer et al., 2023) 210 2 (color, animal obj.)
T2I-CompBench (Huang et al., 2023) 6000 6 (color, counting, texture, shape, (non-)spatial, complex)
Gen-AI (Li et al., 2024a) 1600
8 (scene, attribute, relation, counting,
comparison, differentiation, logic)
ABC-6K (Feng et al., 2023) 6000 1 (color)
WinogroundT2I (Zhu et al., 2023) 22k
20 (action, spatial, direction, color, number, size, texture,
shape, age, weight, manner, sentiment, procedure, speed, etc.)
Comp. Splits (Park et al., 2021) 31k 2 (color, shape)
Winoground (Thrush et al., 2022) 400 5 (object, relation, symbolic, series, pragmatics)
EqBen (Wang et al., 2023) 250k 4 (attribute, location, object, count)
ARO (Yuksekgonul et al., 2023) 50k (relations, attributes)
ConPair (ours) 15k 8 (color, counting, shape, texture, (non-)spatial relations, scene, complex)
Table 1: The comparison of compositional T2I datasets. Contra. is the abbreviation of Contrastive. Complex refers the samples that have multiple objects and complicated attributes and relationships.
2.2 Compositional Datasets and Benchmarks
The most commonly used data sets for object-attribute binding, including DrawBench (Saharia et al., 2022), CC-500 (Feng et al., 2023) and Attend-and-Excite (Chefer et al., 2023) construct text prompts by conjunctions of objects and a few of common attributes like color and shape. To more carefully examine how generative models work on each compositional category, recent work explores the disentanglement of different aspects of text-to-image compositionality. Huang et al. (2023) introduces T2I-CompBench that constructing prompts by LLMs which covers six categories including color, shape, textual, (non-)spatial relationships and complex compositions; Recently, Gen-AI (Li et al., 2024a) collects prompts from professional designers which captures more enhanced reasoning aspects such as differentiation, logic and comparison.
Another line of work proposes contrastive textual benchmarks to evaluate the compositional capability of generative models. ABC-6K (Feng et al., 2023) contains contrast pairs by either swapping the order objects or attributes while they focus on negative text prompts with minimal changes. WinogroundT2I (Zhu et al., 2023) contains 11K complex, high-quality contrastive sentence pairs spanning 20 categories. However, such benchmarks focus on text perturbations but do not have images, which have become realistic with the advancement of generative models.
Several benchmarks featuring contrastive image pairs have also been introduced. Compositional Splits C-CUB and C-Flowers (Park et al., 2021) mainly focused on the color and shape attributes of birds and flowers, sourcing from Caltech-UCSD Birds (Wah et al., 2011), Oxford-102 (Flowers) (Nilsback & Zisserman, 2008). Thrush et al. (2022) curated Winoground consists of 400 high-quality contrastive text-image examples. EqBen (Wang et al., 2023) is an early effort to use Stable Diffusion to synthesize images to evaluate the equivariance of VLMs similarity, but it lacks more complex scenarios. Yuksekgonul et al. (2023) emphasizes the importance of hard negative samples and constructs negative text prompts in ARO by swapping different linguistic elements in the captions sourced from COCO and sampling negative images by the nearest-neighbor algorithm. However, it is not guaranteed the negative images found in the datasets truly match the semantic meaning of the prompts.
Category Stage-I Stage-II
[…]
A red backpack A yellow purse and a blue backpack.
[…]
Two birds. Two dogs and one cat.
[…]
A fluffy toy. A rubber mirror and a glass tire
[…]
A bee on the right of a place.
[…]
A basketball player is dancing. A man is passing a tennis ball to a woman.
[…]
A rainy night. In a serene lake on a sunny day.
[…]
Three square clock. A brown dog are eating pears to the left of two fluffy cats.
Stage-III
Complex Two green birds standing next to two orange birds on a willow tree.
An orange bird standing next to three green birds on the grass.
A man wearing a blue hat is throwing an american football from the left to the right
to a woman wearing a green pants on the playground during a snowy day.
A woman wearing a green hat is throwing a tennis ball from the right to the left
to a woman wearing a blue hat on the playground during a rainy night.
Table 2: Examples of text prompts. Each sample has a positive (top) and a negative prompt (bottom).
3 Data Construction: ConPair
To address attribute binding and compositional generation, we propose a new high-quality contrastive dataset, ConPair. Next, we introduce our design principle for constructing ConPair. Each sample in ConPair consists of a pair of images ($x^{+},x^{-}$) associated with a positive caption $t^{+}$.
3.1 Generating text prompts
Our text prompts cover eight categories of compositionality: color, shape, texture, counting, spatial relationship, non-spatial relationship, scene, and complex. To obtain prompts, we utilize the in-context learning capability of LLMs. We provide hand-crafted seed prompts as examples and predefined templates (e.g., “A {color} {object} and a {color} {object}.”) and then ask GPT-4 to generate similar textual prompts. We include additional instructions that specify the prompt length, no repetition, etc. In total, we generate 15400 positive text prompts. More information on the text prompt generation is provided in the appendix A.
To generate a negative text prompt $t^{-}$, we use GPT-4 to perturb the specified attributes or relationships of the objects for Stage-I data. In Stage-II, we either swap the objects or the attributes, depending on which option makes more sense in the given context. For complex sentences, we prompt GPT-4 to construct contrastive samples by altering the attributes or relationships within the sentences. subsection 2.2 presents our example contrastive text prompts.
3.2 Generating contrastive images
Minimal Visual Differences.
Our key idea is to generate contrastive images that are minimally different in visual representations. By ”minimal,” we mean that, aside from the altered attribute/relation, other elements in the images remain consistent or similar. In practice, we source negative image samples in two ways: 1) generate negative images by prompting negative prompts to diffusion models; 2) edit the positive image by providing instructions (e.g., change motorcycle color to red) using MagicBrush (Zhang et al., 2024), as shown at the left of Figure 2.
Figure 2: EvoGen Framework. Data generation pipeline (left) and curricumlum contrastive learning (right). Quality control of image geneartion (bottom): Given a prompt, SD3 generates multiple candidate images, which are evaluated by LLaVA. We select the best image by alignment and CLIPScore. If the alignment score is low, we prompt LLaVA to describe the image as a new revised caption based on the generated image.
Text-Image Alignment.
The high-level objective of ConPair is to generate positive images that faithfully adhere to the positive text guidance, while the corresponding negative images do not align with the positive text, despite having minimal visual differences from the positive images. As the quality of images generated by diffusion-based T2I generative models varies significantly (Karthik et al., 2023), we first generate 10-20 candidate images per prompt. However, how to select the most faithful image is difficult. Existing automatic metrics like CLIPScore are not always effective at comparing the faithfulness of images when they are visually similar. To address this, we propose decomposing each text prompt into a set of questions using an LLM and leverage the capabilities of VQA models to rank candidate images by their alignment score, as illustrated in Figure 2 (bottom)  [^1] Examples of decomposed questions are provided in the Appendix A.3. Note the correct answers can be directly extracted from the prompts. Intuitively, we consider an image a success if all the answers are correct or if the alignment is greater than $\theta_{\text{align}}$ for certain categories, such as Complex. After getting aligned images, we select the best image by automatic metric (e.g., CLIPScore).
Empirically, we find this procedure fails to generate faithful images particularly when the prompts become complex, as limited by the compositionality understanding of existing generative models, which aligns with the observations of Sun et al. (2023). In response to such cases–i.e., the alignment scores for all candidate images are low–we introduce an innovative reverse-alignment strategy. Instead of simply discarding low-alignment images, we leverage a VLM to dynamically revise the text prompts based on the content of the generated images. By doing so, we generate new captions that correct the previous inaccuracies while preserving the original descriptions, thereby improving the alignment between the text and image.
Image-Image Similarity.
Given each positive sample, we generate 20 negative images and select the one with the highest similarity to the corresponding positive image, ensuring that the changes between the positive and negative image pairs are minimal. In case of color and texture, we use image editing rather than generation, as it delivers better performance for these attributes. Han et al. (2024b) proposes that human feedback plays a vital role in enhancing model performance. For quality assurance, 3 annotators randomly manually reviewed the pairs in the dataset and filtered 647 pairs that were obviously invalid.
Figure 3: Contrastive dataset examples. Each pair includes a positive image generated from the given prompt (left) and a negative image that is semantically inconsistent with the prompt (right), differing only minimally from the positive image.
4 EvoGen: Curriculum Contrastive Fine-tuning
A common challenge in training models with data of mixed difficulty is that it can overwhelm the model and lead to suboptimal learning (Bengio et al., 2009). Therefore, we divide the dataset into three stages and introduce a simple but effective multi-stage fine-tuning paradigm, allowing the model to gradually progress from simpler compositional tasks to more complex ones.
Stage-I: Single object.
In the first stage, the samples consist of a single object with either a specific attribute (e.g., shape, color, quantity, or texture), a specific action, or within a simple static scene. The differences between the corresponding negative and positive images are designed to be clear and noticeable. For instance, “A man is walking” vs. “A man is eating”, where the actions differ significantly, allowing the model to easily learn to distinguish between them.
Stage-II: Object compositions.
We compose two objects with specified interactions and spatial relationships. An example of non-spatial relationship is “A woman chases a dog” vs. “A yellow dog chases a woman.” This setup helps the models learn to differentiate the relationships between two objects.
Stage-III: Complex compositions.
To further complicate the scenarios, we propose prompts with complex compositions of attributes, objects, and scenes. Data in this stage can be: 1) contain more than two objects; 2) assign more than two attributes to each object, or 3) involve intricate relationships between objects.
Ultimately, our goal is to equip the model with the capability to inherently tackle challenges in compositional generation. Next, we discuss how to design the contrastive loss during fine-tuning at each stage. Given a positive text prompt $t$, a generated positive image $x^{+}$, and corresponding negative image $x^{-}$, the framework comprises the following three major components:
Diffusion Model.
The autoencoder converts the positive image and negative image to latent space as $z_{0}^{+}$ and $z_{0}^{-}$. The noisy latent at timestep $t$ is represented as $z_{t}^{+}$ and $z_{t}^{-}$. The encoder of the noise estimator $\epsilon_{\theta}$ is used to extract feature maps $z_{et}^{+}$ and $z_{et}^{-}$ respectively.
Projection head.
We apply a small neural network projection head $g(\cdot)$ that maps image representations to the space where contrastive loss is applied. We use a MLP with one hidden layer to obtain $h_{t}=g(z_{et})=W^{(2)}\sigma(W^{(1)}(z_{et}))$.
Contrastive loss.
For the contrastive objective, we utilize a variant of the InfoNCE loss (van den Oord et al., 2019), which is widely used in contrastive learning frameworks. This loss function is designed to maximize the similarity between the positive image and its corresponding text prompt, while minimizing the similarity between the negative image and the same text prompt. The loss for a positive-negative image pair is expressed as follows:
$\mathcal{L}=-\log\frac{\exp({\text{sim}(h_{t}^{+},f(t))}/{\tau})}{\exp({\text{sim}(h_{t}^{+},f(t))}/{\tau})+\exp({\text{sim}(h_{t}^{-},f(t))}/{\tau})}$ (2)
where $\tau$ is a temperature parameter, $f(\cdot)$ is CLIP text encoder, $\mathrm{sim}$ function represents cosine similarity:
$\mathrm{sim}(u,v)=\frac{u^{T}\cdot v}{\|u\|\|v\|}$ (3)
This encourages the model to distinguish between positive and negative image-text pairs.
5 Experiments and Discussions
5.1 Implementation Details
Model Attribute Binding Object Relationship Complex
Color Shape Texture Spatial Non-Spatial
SD v1.4 (Rombach et al., 2022) 37.65 35.76 41.56 12.46 30.79 30.80
DALL-E 2 (Ramesh et al., 2022) 57.00 55.00 63.74 13.00 30.00 37.00
SDXL (Podell et al., 2023) 64.00 54.00 36.45 20.00 31.00 41.00
Composable v2 (Liu et al., 2023) 40.63 32.99 36.45 8.00 29.80 28.98
Structured v2 (Feng et al., 2023) 49.90 42.18 49.00 13.86 31.11 33.55
Attn-Exct v2 Chefer et al. (2023) 64.00 45.17 59.63 14.55 31.09 34.01
GORs (Huang et al., 2023) 66.03 47.85 62.87 18.15 31.93 33.28
PixArt-$\alpha$ (Chen et al., 2023) 68.86 55.82 70.44 20.82 31.79 41.17
MARS (He et al., 2024) 69.13 54.31 71.23 19.24 32.10 40.49
DALL-E 3 (Betker et al., 2023) 77.85 62.05 70.36 28.65 30.03 37.73
SD v2.1 (Rombach et al., 2022) 50.65 42.21 49.22 13.42 30.96 33.86
EvoGen-SD v2.1 (Ours) 71.04 54.57 72.34 21.76 33.08 42.52
SD3-Medium (Esser et al., 2024) 80.63 58.48 75.24 34.88 31.17 47.72
EvoGen- SD3-Medium (Ours) 83.08 62.83 77.59 36.95 33.12 49.07
Table 3: Alignment evaluation on T2I-CompBench. The best results are in bold
Basic Advanced
Method Attribute Scene Relation Avg Count Differ Compare Logical Avg
Spatial Action Part Negate Universal
SD v2.1 0.75 0.77 0.72 0.72 0.69 0.74 0.66 0.63 0.61 0.50 0.57 0.58
SD-XL Turbo 0.79 0.82 0.77 0.78 0.76 0.79 0.69 0.65 0.64 0.51 0.57 0.60
DeepFloyd-IF 0.82 0.83 0.80 0.81 0.80 0.81 0.69 0.66 0.65 0.48 0.57 0.60
SD-XL 0.82 0.84 0.80 0.81 0.81 0.82 0.71 0.67 0.64 0.49 0.57 0.60
Midjourney v6 0.85 0.88 0.86 0.86 0.85 0.85 0.75 0.73 0.70 0.49 0.64 0.65
SD3-Medium 0.86 0.86 0.87 0.86 0.88 0.86 0.74 0.77 0.72 0.50 0.73 0.68
DALL-E 3 0.91 0.91 0.89 0.89 0.91 0.90 0.78 0.76 0.70 0.46 0.65 0.65
EvoGen- SD3-Medium (Ours) 0.89 0.88 0.90 0.91 0.88 0.89 0.80 0.79 0.73 0.51 0.73 0.72
Table 4: Gen-AI Benchmark Results.
Figure 4: Average CLIP image-text similarities between the text prompts and the images generated by different models. The Full Prompt Similarity considers full text prompt. Minimum Object represents the minimum of the similarities between the generated image and each of the two object prompts. Example of this benchmark is in Appendix C.
Experimental Setup
In an attempt to evaluate the faithfulness of generated images, we use GPT-4 to decompose a text prompt into a pair of questions and answers, which serve as the input of our VQA model, LLaVA v1.5 (Liu et al., 2024). Following previous work (Huang et al., 2023; Feng et al., 2023), we finetune and evaluate EvoGen on SD v2.1 (Rombach et al., 2022) and SD3-Medium (Esser et al., 2024).
Baselines
We compare our results with several state-ofthe-art methods, including trending open-sourced T2I models that trained on large training data, SD v1.4, SD v2.1 and SD3-Medium, DALL-E 2 (Ramesh et al., 2022) and SDXL (Podell et al., 2023). ComposableDiffusion v2 (Liu et al., 2023) is designed for conjunction and negation of concepts for pretrained diffusion models. StructureDiffusion v2 (Feng et al., 2023), Divide-Bind (Li et al., 2024b) and Attn-Exct v2 (Chefer et al., 2023) are designed for attribute binding for pretrained diffusion models. GORs (Huang et al., 2023) finetunes SD v2.1 with selected samples and rewards. PixArt-$\alpha$ (Chen et al., 2023) incorporates cross-attention modules into the Diffusion Transformer. MARS (He et al., 2024) adapts from auto-regressive pre-trained LLMs for T2I generation tasks.
Evaluation Metrics
To quantitatively assess the efficacy of our approach, we comprehensively evaluate our method via three primary metrics: 1) compositionality on T2I-CompBench (Huang et al., 2023), 2) GenAI-Bench Li et al. (2024a) and 3) color-object compositionality prompts (Chefer et al., 2023).
5.2 Performance Comparison and Analysis
Alignment Assessment.
Figure 5: Average CLIP similarity of image-text pairs in ConPair.
To examine the quality of ConPair, we measure the alignment of the positive image and texts using CLIP similarity. Figure 5 compares directly selecting the best image based on CLIPScore with our pipeline, which leverages a VQA model to guide image generation. These results confirm that our approach consistently improves image faithfulness across all categories with VQA assistance during image generation and demonstrate ConPair contains high-quality image-text pairs.
Benchmark Results
Beyond the above evaluation, we also assess the alignment between the generated images using EvoGen and text condition on T2I-Compbench. As shown in Table 3, EvoGen achieves SOTA performance across all categories. In particular, we observe significant improvements in the Non-spatial and Complex categories. We hypothesize this is primarily attributed to Stage-III training, where high-quality contrastive samples with complicated compositional components are leveraged to achieve superior alignment capabilities. We further evaluate EvoGen on the Gen-AI benchmark. As indicated in Table 4, EvoGen performs best on all the Advanced prompts, although it exhibits relatively weaker performance in some of the basic categories compared to DALL-E 3.
Figure 4 presents the average image-text similarity on the benchmark proposed by Chefer et al. (2023), which evaluates the composition of objects, animals, and color attributes. Compared to other diffusion-based models, our method consistently outperforms in both full and minimum similarities across three categories, except for the minimum similarity on Object-Object prompts. These results demonstrate the effectiveness of our approach.
Model Attribute Binding Object Relationship Complex
Color Shape Texture Spatial Non-Spatial
SD v2.1 (Rombach et al., 2022) 50.65 42.21 49.22 13.42 30.96 33.86
ConPair 63.63 47.64 61.64 17.77 31.21 35.02
ConPair + Contra. Loss 69.45 54.39 67.72 20.21 32.09 38.14
ConPair + Contra. Loss + Multi-stage FT (EvoGen-SD v2.1) 71.04 54.57 72.34 21.76 33.08 42.52
Table 5: Ablation on T2I-CompBench. ConPair refers to directly finetune SD v2.1 on ConPair.
Figure 6: Qualitative comparison between EvoGen and other SOTA T2I modes with different prompts.
Ablation Study
We conduct ablation studies on T2I-CompBench by exploring three key design choices. First, we assess the effectiveness of our constructed dataset, ConPair, by fine-tuning SD v2.1 directly using ConPair. As shown in Table 5, our results consistently outperform the baseline evaluation on SD v2.1 across all categories, demonstrating that our data generation pipeline is effective. Next, we validate the impact of our contrastive loss by comparing it with fine-tuning without this loss. The contrastive loss improves performance in the attribute binding category, though it has less impact on object relationships and complex scenes. We hypothesize this is because attribute discrepancies are easier for the model to detect, while relationship differences are more complex. Finally, applying the multi-stage fine-tuning strategy leads to further improvements, particularly in the Complex category, suggesting that building a foundational understanding of simpler cases better equips the model to handle more intricate scenarios.
Qualitative Evaluation
Figure 7: Examples of EvoGen for complex compositionality.
Figure 6 presents a side-by-side comparison between EvoGen and other state-of-the-art T2I models, including SDXL, DALL-E 3, SD v3 and PixArt-$\alpha$. EvoGen consistently outperforms the other models in generating accurate images based on the given prompts. SDXL frequently generates incorrect actions and binds attributes to the wrong objects. DALL-E 3 fails to correctly count objects in two examples and misses attributes in the first case. SD v3 struggles with counting and attribute binding but performs well in generating actions. PixArt-$\alpha$ is unable to handle attributes, spatial relationships, and fails to count objects accurately in the second prompt.
Next, we evaluate how our approach handles complex compositionality, as shown in Figure 7. Using the same object, “bear” and “cat,” we gradually increase the complexity by introducing variations in attributes, counting, scene settings, interactions between objects, and spatial relationships. The generated results indicate that our model effectively mitigates the attribute binding issues present in existing models, demonstrating a significant improvement in maintaining accurate compositional relationships.
Figure 8: User study on 100 randomly selected prompts from Feng et al. (2023). The ratio values indicate the percentages of participants preferring the corresponding model.
User Study
We conducted a user study to complement our evaluation and provide a more intuitive assessment of EvoGen’s performance. Due to the time-intensive nature of user studies involving human evaluators, we selected top-performing comparable models—DALLE-2, SD v3, SDXL, and PixArt-$\alpha$—all accessible through APIs and capable of generating images. As shown in Figure 8, the results demonstrate EvoGen’s superior performance in alignment, though the aesthetic quality may be slightly lower compared to other models.
6 Conclusion
In this work, we present EvoGen, a curriculum contrastive framework to overcome the limitations of diffusion models in compositional text-to-image generation, such as incorrect attribute binding and object relationships. By leveraging a curated dataset of positive-negative image pairs and a multi-stage fine-tuning process, EvoGen progressively improves model performance, particularly in complex scenarios. Our experiments demonstrate the effectiveness of this method, paving the way for more robust and accurate generative models.
7 Limitation
Despite the effectiveness of our current approach, there are a few limitations that can be addressed in future work. First, our dataset, while comprehensive, could be further expanded to cover an even broader range of compositional scenarios and object-attribute relationships. This would enhance the model’s generalization capabilities. Additionally, although we employ a VQA-guided image generation process, there is still room for improvement in ensuring the faithfulness of the generated images to their corresponding prompts, particularly in more complex settings. Refining this process and incorporating more advanced techniques could further boost the alignment between the text and image.
8 Reproducibility
We have made efforts to ensure that our method is reproducible. Appendix A provides a description of how we construct our dataset. Especially, Appndix A.1 and A.2 presents how we prompt GPT-4 and use predefined template to generate text prompts of our dataset. Appendix A.3 provides an example how we utilize VQA system to decompose a prompt into a set of questions, and answers. Appendix B provides the details of implementation, to make sure the fine-tuning is reproducible.
Appendix A ConPair Data Construction
A.1 Text prompts generation
Here, we design the template and rules to generate text prompts by GPT-4 as follows:
  • Color: Current state-of-the-art text-to-image models often confuse the colors of objects when there are multiple objects. Color prompts in Stage-I follow fixed sentence template “A {color} {object}.” and “A {color} {object} and a {color} {object}.” for Stage-II.
  • Texture: Following Huang et al. (2023), we emphasize in the GPT-4 instructions to require valid combinations of an object and a textural attribute. The texture prompts follows the template “A {texture} {object}.” for Stage-I and “A {texture} {object} and a {texture} {object}.” for Stage-II.
  • Shape: We first generate objects with common geometric shapes using fixed template “A {shape} {object}.” for Stage-I and “A {shape} {object} and a {shape} {object}.” for Stage-II. Moreover, we ask GPT-4 to generate objects in the same category but with different shapes, e.g., American football vs. Volleyball, as contrastive samples.
  • Counting: Counting prompts in Stage-I follows fixed sentence template {count} {object}.” and {count} {object} and {count} {object}.” for Stage-II.
  • Spatial Relationship: Given predefined spatial relationship such as next to, on the left, etc, we prompt GPT-4 to generate a sentence in a fixed template as {object} {spatial} {object}.” for Stage-II.
  • Non-spatial Relationship: Non-spatial relationships usually describe the interactions between two objects. We prompt GPT-4 to generate text prompts with non-spatial relationships (e.g., actions) and arbitrary nouns. We guarantee there is only one object in the sentence for Stage-I, and two objects in Stage-II. We also find generative models fails to understand texts like “A woman is passing a ball to a man”. It’s hard for the model to correctly generate the directions of actions. We specially design prompts like this.
  • Scene: We ask GPT-4 to generate scenes such as weather, place and background. For Stage-I, the scene is simple, less than 5 words (e.g., on a rainy night.); For Stage-II, scenes combine weather and background or location (e.g., in a serene lake during a thunderstorm.).
  • Complex: Here, we refer to prompts that either contain more than two objects or assign more than two attributes to each object, or involve intricate relationships between objects. We first manually curate 10 such complex prompts, each involving multiple objects bound to various attributes. These manually generated prompts serve as a context for GPT-4 to generate additional natural prompts that emphasize compositionality. The complex cases in Stage-II will be two objects with more attributes; Stage-III involves more objects.
A.2 Negative Text Prompts Generation
We apply in-context learning, everytim we prompt GPT-4 to generate negative cases, we give 5-10 example test prompts each time, and make sure the generation is not repetitive, within certain lengths.
  • In Stage-I, we prompt GPT-4 to change the attribute of the object such as color, shape, texture, counting, action, or scene, with instruction the differences should be noticeable.
  • In Stage-II, we either swap the objects or attributes and let GPT-4 to validate the swapped text prompts. For complex cases, we generate negative text by asking GPT-4 to change the attributes/relationship/scenes.
  • In Stage-III, we carefully curate complicated examples with 3-6 objects, each object have 1-3 attributes, with negative prompts change attributes, actions and spatial relationships, scenes. And we prompt GPT-4 with such examples.
A.3 VQA Assistance
Instruction for QA Generation. Given an image description, generate one or two multiple-choice questions that verify if the image description is correct. Table 6 shows an example of a generated prompt and QA.
Prompt Question Answer
A brown bear and a white cat, both wearing spacesuits, are playing frisbee on Mars Is there a bear? Yes
Is there a cat? Yes
What color is the bear? Brown
What color is the cat? White
Does the bear wear a spacesuit? Yes
Does the cat wear a spacesuit? Yes
Is the bear playing the frisbee? Yes
Is the cat playing the frisbee? Yes
Where are they playing? Mars
Table 6: VQA generated questions from a prompt.
Next, we illustrate how we prompt VQA to revise the caption when alignment scores of all candidate images are low. Given a generated image and a original text prompt, we prompt LLaVA “Given the original text prompt describing the image, identify any parts that inaccurately reflect the image. Then, generate a revised text prompt with correct descriptions, making minimal semantic changes.”. At the same time, we will give example of revised caption pairs to VQA for in-context learning.
A.4 Data Statistics
Stage-I Stage-II Stage-III Total
Shape 500 1000 200 1700
Color 800 1000 200 2000
Counting 800 1000 200 2000
Texture 800 1000 200 2000
Spatial - 1000 200 1200
Non-spatial 800 1000 200 2000
Scene 800 1000 200 2000
Complex - 500 2000 2500
Table 7: Corpus Statistics.
The dataset is organized into three stages, each progressively increasing in complexity. In Stage-I, the dataset includes simpler tasks such as Shape (500 samples), Color (800), Counting (800), Texture (800), Non-spatial relationships (800), and Scene (800), totaling 4,500 samples. Stage-II introduces more complex compositions, with each category—including Shape, Color, Counting, Texture, Spatial relationships, Non-spatial relationships, and Scene—containing 1,000 samples, for a total of 7,500 samples. Stage-III represents the most complex scenarios, with fewer but more intricate samples. We also include some simple cases like Stage-I and II, each contain 200 samples, while the Complex category includes 2,000 samples, totaling 3,400 samples. Across all stages, the dataset contains 15,400 samples, providing a wide range of compositional tasks for model training and evaluation. Figure 9 show more examples of images in our dataset.
Figure 9: Contrastive Images in ConPair
Appendix B Training Implementation Details
We implement our apprpoach upon stable Diffuion v2.1 and we employ the pre-trained text encoder from the CLIP ViT-L/14 model. The VAE encoder is frozen during training. The resolution is 768, the batch size is 16, and the learning rate is 3e-5 with linear decay.
Appendix C Attn & Exct Benchmark Prompt Examples
Model Animal-Animal Animal-Obj Obj-Obj
Stable v1.4 (Rombach et al., 2022) 0.76 0.78 0.77
Composable v2 (Liu et al., 2023) 0.69 0.77 0.76
Structured v2 (Feng et al., 2023) 0.76 0.78 0.76
Attn-Exct v2 (Chefer et al., 2023) 0.80 0.83 0.81
CONFORM (Meral et al., 2023) 0.82 0.85 0.82
Ours 0.84 0.86 0.85
Table 8: Attn-Exct benchmark Results.
The benchmark protocol we follow comprises structured prompts ‘a [animalA] and a [animalB]’, ‘a [animal] and a [color][object]’, ‘a [colorA][objectA] and a [colorB][objectB]’ . Table 8 demonstrate the results of average CLIP similarities between text prompts and captions generated by BLIP for Stable Diffusion-based methods on this benchmark. EvoGen outperform those models on three categories.
Appendix D Qualitative Results
Figure 10 presents more comparison between EvoGen and other state-of-the-art T2I models, including SDXL, DALL-E 3, SD v3 and PixArt-$\alpha$.
Figure 10: Qualitative Results.